OpenStudy (anonymous):

What is the slope of the line which passes through (0, 3) and (0, 5)? –2 0 2 undefined

OpenStudy (btaylor):

slope = (change in x)/(change in y) = (0)/(2) = 0

Parth (parthkohli):

If the \(x\) coordinates remain the same, then the slope is UNDEFINED.

Parth (parthkohli):

@BTaylor Heh, slope = change in y/change in x :)

OpenStudy (btaylor):

my bad. i fail OS algebra 101.

Parth (parthkohli):

lol

OpenStudy (anonymous):

thank you soo much!<3